a meeting of envoys was attended by 4 americans and 5 filipinos. If 6 envoys were selected at random one after the other, determine the values of the random variable F representing the number of filipinos


1
Expert's answer
2022-05-09T18:31:46-0400

The possible values of the random variable F representing the number of filipinos are 2, 3, 4 ,5.


"P(F=2)=\dfrac{\dbinom{4}{4}\dbinom{5}{2}}{\dbinom{4+5}{6}}=\dfrac{10}{84}"

"P(F=3)=\dfrac{\dbinom{4}{3}\dbinom{5}{3}}{\dbinom{4+5}{6}}=\dfrac{40}{84}"

"P(F=4)=\dfrac{\dbinom{4}{2}\dbinom{5}{4}}{\dbinom{4+5}{6}}=\dfrac{30}{84}"

"P(F=5)=\dfrac{\dbinom{4}{1}\dbinom{5}{5}}{\dbinom{4+5}{6}}=\dfrac{4}{84}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 f & 2 & 3 & 4 & 5 \\ \hline
 p(f) & 10/84 & 40/84 & 30/84 & 4/84 & f \\
\end{array}"